OpenStudy (anonymous):

Simplify. –24 – 8 ÷ (–4) A. 8 B. –8 C. –22 D. –26

TheSmartOne (thesmartone):

First divide by -4 so -24-8/-4 -24- (-2) =-24+2 =-22

OpenStudy (anonymous):

-22

OpenStudy (anonymous):

tysm aka (thank you so much)